The length of time a full length movie runs from opening to credits is normally distributed with a mean of 1.9 hours and standard deviation of 0.3 hours. Calculate the following :
a. A random movie is between 1.8 and 2.0 hours
b. A movie is longer than 2.3 hours
c. The length of movie that is shorter than 94% of the movies
1
Expert's answer
2020-06-04T20:00:23-0400

Let X = the random variable denoting the length of time a full movie runs from opening to credits


We are given X ~ N("\mu=1.9,\sigma^2=0.3^2)"


Then we have,


Z = "\frac{X-\mu}{\sigma}" ~ N(0, 1), Z is the standard normal variate


a. The probability that a random movie is between 1.8 and 2.0 hours


= P(1.8 "\leq" X "\leq" 2.0)


= P("\frac{1.8-1.9}{0.3}\leq\frac{X-1.9}{0.3}\leq\frac{2.0-1.9}{0.3})"


= P(- 0.33 "\leq" Z "\leq" 0.33)


= "\Phi(0.33)-\Phi(-0.33)"


= 0.6293 - 0.3707


= 0.2586


Answer: The probability that a random movie is between 1.8 and 2.0 hours is 0.2586.


b. The probability that a random movie is longer than 2.3 hours


= P(X > 2.3)


= P("\frac{X-1.9}{0.3}" > "\frac{2.3-1.9}{0.3}")


= P(Z > 1.33)


= 1 - P(Z "\leq" 1.33)


= 1 - "\Phi"(1.33)


= 1 - 0.9082


= 0.0918


Answer: The probability that a random movie is longer than 2.3 hours is 0.0918.


c. Let the length of movie that is shorter than 94% of the movies be t.


Then P(X "\geq" t) = 94%


i.e. P("\frac{X-1.9}{0.3}\geq\frac{t-1.9}{0.3}") = 0.94


i.e. P(Z "\geq\frac{t-1.9}{0.3}") = 0.94


i.e. 1 - P(Z < "\frac{t-1.9}{0.3}") = 0.94


i.e. P(Z < "\frac{t-1.9}{0.3}") = 1 - 0.94 = 0.06


i.e. "\Phi(\frac{t-1.9}{0.3})" = 0.06 = "\Phi("-1.56)


i.e. "\frac{t-1.9}{0.3}" = -1.56


i.e. t = 1.432 = 1.4 (rounded to 1 decimal place)


Answer: The length of movie that is shorter than 94% of the movies is 1.4 hours.